OpenStudy (anonymous):

log 10^3 + log 10^-7

OpenStudy (anonymous):

is this \[\log(10^3)+\log(10^{-7})\]?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

if you don't see the base of the log written, then you can assume it is base ten that makes \(\log(10^3)=3\) and \(\log(10^{-7})=-7\) and so you have the easy task of computing \[3-7\]

OpenStudy (anonymous):

you okay with that?

OpenStudy (anonymous):

is it 4 or -4?

OpenStudy (anonymous):

\(3-7=-4\)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw